Ok so if you got an 88% in your first test, and you got a 69% in your second test, what's your overall average score now?
Answer by stanbon(75887)

Ok so if you got an 88% in your first test, and you got a 69% in your second test, what's your overall average score now?
------
Average = (88+69)/2 = 157/2 = 78.5%
---------------------
Cheers,
Stan H.
--------------